OpenStudy (anonymous):

A vehicle can accelerate or decelerate at a maximum value of 1m/s^2 and can attain a maximum speed of 20m/s if it starts from rest what is the shorest time it can travel 1000m in a staright path if it comes tp rest at the end of 1000m

OpenStudy (anonymous):

@ajprincess we have to find t3 tooo but how???

OpenStudy (amistre64):

t3 = t1 doesnt it?

OpenStudy (anonymous):

No

OpenStudy (amistre64):



OpenStudy (anonymous):

i think we have to find the distance but sont knw hw??

OpenStudy (amistre64):

a(t) = 1 v(t) = t how long does it take to get from 0 to 20?

OpenStudy (anonymous):

20s

OpenStudy (amistre64):

we can integrate v(t) from 0 to 20 to find its displacement \[\int t\ dt=t^2/2\] 0 is trivial, so 20^2/2 = 200 meters to start with

OpenStudy (anonymous):

1st it is travellling with acceleration next with constant speed and last with deceleration we have to find the distance

OpenStudy (amistre64):

it slows down as fast as it speeds up, right?

OpenStudy (amistre64):

so t1 = t3

OpenStudy (anonymous):

yes u r correct

OpenStudy (anonymous):

yes t1=t3 we have to find t2 at constant speed

OpenStudy (amistre64):

400 meters is taken up by speeding up and slowing down

OpenStudy (amistre64):

600 meters left to go 20 meters per sec

OpenStudy (amistre64):

20t = 600 when t = 30 i believe